OpenStudy (anonymous):

Choose the letter for the best answer. An isosceles triangle has two sides that are equal in length. Isosceles triangle ABC is similar to isosceles triangle XYZ. What proportion would you use to find the length of the third side of triangle XYZ? A. BC/XZ = AB/XY B. AC/XY = BC/XZ C. AB/XY = AC/XZ D. AB/XY = BC/YZ **Must show work! (Metals)

OpenStudy (anonymous):

did I not get it right?

OpenStudy (anonymous):

The answer is correct but I still don't get it, it would be better if you show the work so I can understand me and it does because it helps me.

OpenStudy (anonymous):

i did show the work...

OpenStudy (anonymous):

It doesn't look like showing work.

OpenStudy (anonymous):

here: 9/6 = 12/x ~ AB/XY = AC/XZ (Original Equation) 72 = 9x (Cross Multiply) 8 = x (Simplify by dividing both sides by 9)

OpenStudy (anonymous):

that work?

OpenStudy (anonymous):

Hmm, I guess. But I kinda understand.. but can you draw it out?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):